# Real-Valued Somewhat-Pseudorandom Unitaries

### v1

@@ 4.3 Proving the main lemma and theorem
Let $s,t$ be polynomials in $n$ and let $\{|{\psi^{(j)}}\rangle\}_{j\in[s]}$ be orthogonal quantum states. From corollary 4.4, choosing $c=8$ we get that $\forall_{j\in[s]}H^{\otimes n}U_{f}|{\psi^{(j)}}\rangle$ is $\frac{8n}{N}$-flat with probability at least $1-s\cdot 2\exp\left(-\left(\frac{8}{4}-\ln(2)\right)n\right)\geq 1-2^{-n}=1-\frac{1}{N}$. Assuming flatness of these states, we can now use lemma 4.5 and get that:
@@
Taking $Gen_{n}(k)=U_{k}$ to be $U_{F,G,P}$ (where $k$ is split into keys for $F,G$ and $P$), we get that it is indeed a QPT algorithm on $n$ qubits. We now prove the security requirement. Recall that $\rho_{in}$ is promised to be of the form
@@
$\left\|\Phi(\rho_{in})-\sum_{a\in A}p_{a}(\rho_{a}\otimes\rho_{\text{\sf uni}_{s,t}})\right\|\leq O\left(\frac{s^{6}t^{4}n^{2}}{N}\right)~{}.$ (113)
@@
We finish by combining equations 115 and 114 to get
@@
as needed to satisfy the security definition 3.1. ∎

### v2

@@ 4.3 Proving the main lemma and theorem
Let $s,t$ be polynomials in $n$ and let $\{|{\psi^{(j)}}\rangle\}_{j\in[s]}$ be orthogonal quantum states. From corollary 4.4, choosing $c=8$ we get that $\forall_{j\in[s]}H^{\otimes n}U_{f}|{\psi^{(j)}}\rangle$ is $\frac{8n}{N}$-flat with probability at least $1-s\cdot 2\exp\left(-\left(\frac{8}{4}-\ln(2)\right)n\right)\geq 1-2^{-n}=1-\frac{1}{N}$.
Assuming flatness of these states, we can now use lemma 4.5 and get that:
@@
Taking $Gen_{n}(k)=U_{k}$ to be $U_{F,G,P}$ (where $k$ is split into keys for $F,G$ and $P$), we get that it is indeed a QPT algorithm on $n$ qubits. We now prove the security requirement.
Recall that $\rho_{in}$ is promised to be of the form
@@
$\left\|\Phi(\rho_{in})-\sum_{a\in A}p_{a}(\rho_{a}\otimes\rho_{\text{\sf uni}_{s,t}})\right\|_{1}\leq O\left(\frac{s^{6}t^{4}n^{2}}{N}\right)~{}.$ (87)
@@
We finish by combining equations 89 and 88 to get
@@
as needed to satisfy the security definition 3.1.